Moody wants to find the height of the tallest building in his city. He stands 332 feet away from the building. There is a tree 34 feet in front of him, which he knows is 20 feet tall. How tall is the building to the nearest foot?
Answer by checkley77(12844)

Assuming he is lying on the ground then:
20/34=x/332
34x=20*332
34x=6,640
x=6,640/34
x=195 ft. tall.